Two equally charged particles are held 3.2×10-3 m apart and then released from rest. The initial acceleration of the first particle is observed to be 7.0 m/s2 and that of the second to be 9.0 m/s2. If the mass of the first particle is 6.3 x l0-7 kg, what are (a) the mass of the second particle and (b) the magnitude of the charge of each particle


1
Expert's answer
2022-04-14T15:46:27-0400

"\frac{a_1}{a_2}=\frac{m_1}{m_2},"

"m_2=m_1\frac{a_2}{a_1},"

"m_2=8.1\cdot 10^{-7}~kg."